Ripples in a pond travel at a velocity of 3 m/s with one peak passing a reed near the shore every 0.25 seconds. What is the wavelength of the ripples? What is their frequency?
Answer #1

speed, v = 3m/s

time period = time between two peaks, T= 0.25 sec

now wavelength = v ×T = 3 ×0.25 = 0.75 m

frequency, f =1/T = 1/0.25 = 4 Hz